# Learning Coherent Matrixized Representation in Latent Space for Volumetric 4D Generation

### v2

@@ 11 Visualization of Unconditional 4D Sequence Shape Generation
12 Nearest Sampling
We also show the most similar objects in DT4D dataset to unconditional generation results in Fig.18.
@@ 12 Nearest Sampling
13 Discussion of 4D Generative Model
We compare L4GM and 4Diffusion (video-based) using LPIPS and CLIP-S for overall quality, CLIP-T and FVD for consistency (30s), AfC for artifacts, and measure max action length and processing time for efficiency in Tab. 12.
Tab. 12 indicates that L4GM’s 3D Gaussian representation excels in appearance quality (LPIPS, CLIP) but is computationally expensive with limited sequence length. Video-based methods offer visual richness but lack geometric guidance, leading to illusions. Our method outperforms in efficiency (time), consistency (CLIP-T, FVD), sequence length, and avoids illusions, highlighting the superiority of our volume-based design.
14 Quality at different scale, Consistency and Alignment
We conduct more experiments on the 110k subset and 323k of Objaverse in Tab. 13 to validate the quality at different scales, spatiotemporal consistency, and alignment with conditions through a user study, where participants are asked to rate each example in these three aspects on a scale from 1 to 5. Our model outperforms STAG4D on all metrics.
@@ 14 Quality at different scale, Consistency and Alignment
Table 13: User study ratings for generated 4D sequences on 110k and 323k subsets of Objaverse.
Method Average CLIP-S
Ours 0.903
STAG4D 0.896
Table 14: Model performance on in-the-wild inputs.
Figure 19: Wild input conditional 4D generation results.
15 Evaluation on Wild Input
We also test model’s ability on 20 in-the-wild Internet inputs in Tab. 14, Fig.19 shows our visualization. Compared to STAG4D, our model demonstrates stronger generalization ability and more stable performance.
Figure 20: More conditional 4D generation results on Objaverse.
16 Additional Qualitative Results
We show more conditional volumetric 4D sequence generation examples in Fig 21 (Page.5) and  20.
Figure 21: More conditional 4D generation results on 3DBiCar.

### v3

@@ 4.2 Unconditional 4D Sequence Shape Generation
Fig. 8 explores motion diversity and interpolation, with each row depicting a different action performed by the same object. This demonstrates our capability to generate varied motions for the same object with high quality, highlighting both its versatility and its detail-rich output. Additionally, the 2nd, 4th, and 6th columns display objects decoded from interpolated $\mathcal{M}$. A comparison of these columns with their adjacent ones reveals that the interpolated results exhibit strong temporal consistency, indicating that our latent matrixized 4D representation supports the generation of sequences with variable-length frame density through interpolation. Because our matrixized 4D representation forms a continuous and meaningful latent space after sufficient training, enabling the construction of variable-length 4D sequences through interpolation.
4.3 Conditonal 4D Sequence Generation
Figure 11: Generated volumetric 4D sequence.
Figure 12: Identity transfer experiments.
Method Speed Color CLIP-S
3DBiCar Obj.-110k Obj.-323k
DG4D [29] 6.5 min $\checkmark$ 0.803 0.790 0.785
Hyper. [9] 0.54s $\times$ - - -
STAG4D [38] 7.5 min $\checkmark$ 0.908 0.895 0.906
L4GM [30] 3.7s $\checkmark$ 0.940 0.925 0.930
GenMOJO [6] 46.3s $\checkmark$ 0.909 0.911 0.903
Ours 0.625s $\checkmark$ 0.911 0.905 0.915
Figure 13: Comparison of 4D generation methods.
Generation via conditions. First, we present our conditional 4D sequence generation results in Fig. 9. We show images from different viewing angles and depth maps of the generated sequences to demonstrate the ability to produce plausible, topologically continuous, and visually consistent 4D sequences. Fig. 9 illustrates that our model effectively learns color and shape information, and generates 4D motion sequences from image and text prompts. Because our method generates the 4D representation $\mathcal{M}$ by balancing the trade-off between 3D shape quality and sequence coherence, which can be accurately decoded into 4D sequences maintaining diversity and alignment with image-text guidance, benefiting from our coherent 3D shape and color modeling. Meanwhile, benefiting from our HCSTA attention module, we explicitly supervise the correlation between shape and color while maintaining spatiotemporal consistency.
Variable-length 4D sequences. To demonstrate the capability of generating arbitrarily long 4D sequences, we create latent 4D representations of the same object performing different actions using varying motion prompts, then apply interpolation to generate intermediate frames. Fig. 10 shows that our method can produce variable-length 4D sequences through identity transfer and interpolation. This validates the practicality of our approach for generating coherent 4D sequences in stages and interpolations. Because after sufficient training, our matrixized 4D representation forms a continuous and meaningful latent space, allowing variable-length 4D sequences to be constructed via interpolation and our coherent 3D shape and color modeling.
Volumetric property. To demonstrate the volumetric property of our generated results, we create volumetric 4D sequences, place them in a shared scene, and perform free navigation and rendering. Results in Fig. 11 show that our method supports navigation across different views (a), (b), and (c) while maintaining high-quality rendering. Additionally, our approach enables free editing and manipulation of each frame, distinguishing it from other free-form 4D generation models. This is because our coherent 3D shape and color modeling, together with the matrixized 4D representation, constructs volumetric 4D sequences in which each frame is a high-fidelity, editable mesh. Compared to 4D representations based on 3D Gaussian Splatting or video generation, our method offers superior editability and better compatibility with standard computer graphics pipelines, distinguishing it from existing approaches.
Cross Identity transfer. To evaluate the controllability of our method, we apply the same motion prompt to images of different identities, generating 4D sequences with consistent motion but varying identities. Results in Fig. 13 show that our method retains the input’s identity and appearance while accurately capturing the motion described in the prompt. This is because our latent matrixized 4D sequence representation effectively decouples shape, color, and motion. With the HCSTA layer, we can explicitly supervise the spatiotemporal information of shape, appearance, and motion, enabling the same motion to be performed across different objects.
Comparison of 4D Generative Models. We first compare overall quality and generation speed. Our model uses diffusion to generate latent 4D sequence representation $\mathcal{M}$ from image and text conditions. First, we compare it with DreamGaussian4D (DG4D) [29], HyperDiffusion [9], STAG4D [38], L4GM [30], and GenMOJO [6] in Tab. 13 using 3DBiCar and Objaverse 110k and 323k subsets. DG4D supports colored 4D sequences but is slow and memory-intensive. Although STAG4D significantly enhances output quality, this improvement is accompanied by increased computational overhead. HyperDiffusion, while fast, compromises on both color and geometric quality. Although L4GM and 4Diffusion achieve similar generation quality to ours, they require significantly more time and memory. In contrast, our method balances speed and memory efficiency in generation by leveraging matrixized 4D representation.
We also test the spatiotemporal consistency and the alignment with the input using L4GM and 4Diffusion (video-based) using LPIPS and CLIP-S for overall quality, CLIP-T and FVD for consistency (30s), AfC for the average number of artifacts, and measure max action length and processing time for efficiency in Tab. 3.
@@ 4.3 Conditonal 4D Sequence Generation
[…]
figureQualitative comparison with DreamGaussian4D.
@@
[…]
tableUser study ratings for generated 4D sequences on 110k and 323k subsets of Objaverse.
Tab. 3 indicates that L4GM’s 3D Gaussian representation excels in appearance quality (LPIPS, CLIP) but is computationally expensive with limited sequence length. Video-based methods offer visual richness but lack geometric guidance, leading to illusions. Our method outperforms in efficiency (time), consistency (CLIP-T, FVD), sequence length, and avoids illusions. This is because our volumetric 4D design explicitly incorporates geometric information and decouples shape from color, enabling explicit supervision of per-frame geometry and appearance. This significantly enhances spatiotemporal consistency and reduces the average number of artifacts.
We conduct more experiments on the 110k subset and 323k of Objaverse in Tab. 4.3 to validate the quality at different scales, spatiotemporal consistency, and alignment with conditions through a user study, where participants are asked to rate each example in these three aspects on a scale from 1 to 5. Our model outperforms STAG4D on all metrics. This is because our method maintains better temporal coherence and input alignment by explicitly decoupling shape and color, and employing a staged HCSTA attention module to enhance input consistency and explicitly supervise spatiotemporal coherence.
We further compare the geometric quality with 4D generation models based on 3D Gaussian Splatting. We compare our method with DG4D in Fig. 4.3, providing color images and normal maps from different views to highlight the differences. Due to GPU limitations, DG4D exhibits poor rendering quality outside supervised views. In contrast, our model generates 4D sequences with more complete appearances and smoother normal and depth maps. Our approach outperforms DG4D by balancing memory efficiency with high-quality sequence generation. This is because our method accurately models per-frame geometry and geometry-aware color through coherent 3D shape and color modeling, focusing on accurate shape representation for each identity, whereas DG4D emphasizes the rendering effects of Gaussian primitives without a precise geometry representation, resulting in lower geometric accuracy.
5 Conclusion
We propose an image-text conditioned volumetric 4D sequence generation method balancing 3D shape quality with sequence coherence, while maintaining diversity and alignment with input guidances. Experiments demonstrate our method’s efficiency and superior performance. We hope it will be a useful attempt that opens up new possibilities for the representation and generation of complex 4D sequences.
Limitations. The lack of large-scale real-world datasets limits 4D generation performance. Improved 3D/4D representations and the use of generative priors will be key areas for future breakthroughs.